Find the 100th term of the arithmetic sequence with first term 9 and eighth term 37


1
Expert's answer
2021-04-25T16:46:47-0400

arithmetic sequence an = a1 + (n-1)*d ; Find d: a8 = a1 + 7*d; d = (a8 - a1)/7 = (37-9)/7 = 28/4=7

Then when n=100, a100 = a1 + (100-1)*d = 9 + 99 * 7 = 702